OpenStudy (anonymous):

Calculate the Lorentz factor for objects trav- eling at 58.9% of the speed of light.

OpenStudy (anonymous):

check me on this, but isn't this the lorentz factor equation? so turn the 58.9% of the speed of light into an actual speed, plug it in for v, c is the speed of light, and calcualte it and ur done ^_^ \[\gamma = \frac{ 1 }{ \sqrt{1 - \frac{ v^2 }{ c^2 }}}\]

OpenStudy (anonymous):

Oh that does make sense :)

OpenStudy (anonymous):

^_^